OpenStudy (anonymous):

find the equation of the derivative function y=0.3x^2-8x+4

Parth (parthkohli):

Oh, you want to find the derivative of this function, right?

Parth (parthkohli):

In other words, \(\large \rm y'\). Do you know how to find the derivative of the function you're given?

OpenStudy (anonymous):

no o:

Parth (parthkohli):

Do you know the power rule?

OpenStudy (anonymous):

nx^n-1

Parth (parthkohli):

Yep... so can you use the power rule here?

OpenStudy (anonymous):

i dont think so

Parth (parthkohli):

Yes, you can.

Parth (parthkohli):

Can I give you an example?

OpenStudy (anonymous):

mhmm

Parth (parthkohli):

For example, you have to find the derivative function of\[\rm \large y = 3x^2 + 7x -4\]For that, you have to use the Power Rule term by term. First, find the derivative of \(\rm 3x^2\), then \(\rm 7x\) and finally \(\rm -4\). The derivative of \(\rm 3x^2 =2\times3x^{2 - 1} = 6x\) The derivative of \(\rm 7x^1=1\times 7x^{1 - 1} = 1 \times 7x^0 = 1 \times 7 \times 1 = 7\) The derivative of \(\rm -4 = 0\) because the derivative of a constant is 0. So, the derivative of this whole thing can be written as\[\rm \large y'=6x +7 -0 = 6x + 7\]Got it?